# Galled Perfect Transfer Networks

### v2

@@ 2.3 Differences with reconstructions from bipartitions
As we pointed out in the introduction, Warnow et al. [warnow2024statistically] also independently developed galled tree reconstructions algorithms from character data (which are assumed to be SNPs in the paper). Therein, for each character, the species can be in one of two states, and the states are assumed to correspond to a bipartition of a galled tree $N$. A bipartition is a partition of the leaves that can be obtained by taking a tree displayed by $N$, removing an edge, and taking the two sets of leaves of the resulting connected components (here, “displayed” means a tree that can be obtained by removing one parent edge from each node of in-degree two). They discuss algorithms to reconstruct a galled tree that contains a set of given characters, that is, that contains all the corresponding bipartitions. More specifically, they show that if all the bipartitions of a galled tree $N$ are known (but $N$ is unknown), then $N$ can be reconstructed in polynomial time. They also show that this approach is statistically consistent if characters evolve on a galled tree and may or may not be transmitted on edges labeled as transfer.
@@ 3 The Galled Completion Problem
In other words, $v$ is a FA for $C$ if it roots a maximal subtree of taxa that contain $C$. An example of this definition is shown in Figure 3.
@@
For a character $C\in\mathcal{C}$ and a network $N$, we will say that a node $v$ is an origin for $C$ if $v$ reaches every leaf in $C$ in $N-F_{C}(N)$. By Lemma 1, our goal is to add arcs to $T$ to ensure that every character has an origin. We first show that in any galled-completion of a tree $T$, an origin of a character $C$ must descend from a FA node (or it could be a transfer node added just above it), and it must “give” its character to all other FAs by transferring just above them.
@@
By definition, in $T$ the ancestors of FAs are in $F_{C}(N)$ since they have descendants not in $C$. Thus, an origin $w$ in $N$ must be in a FA subtree, or it could be a transfer node inserted above such an FA during the process of adding transfer edges from $T$ to $N$. If there are multiple FAs, $w$ must be able to reach the other ones, and Lemma 2.2 lets us establish that transfer edges as described are needed to achieve this. ∎
@@
We focus on the first condition. First observe that if $v\in V(T)$, then $v\in V(N)$ as well. Moreover, the the parent of $v$ in $\mathcal{T}_{N}$ is either also the parent $p_{v}$ of $v$ in $T$, or it could be a transfer node created to complete $T$ into $N$. In the latter case, note that at most one transfer node could could be present on the path between $p_{v}$ and $v$ in $\mathcal{T}_{N}$, as otherwise, there would be two consecutive transfer nodes which, as one can check, would violate Lemma 2.2 ($p_{v}$ would have two descendants with outgoing transfers). This lets us establish that some edges of $T$ are subdivided to obtain $\mathcal{T}_{N}$, but each edge can be subdivided at most once.
@@
Let us next observe that if $w$ is an origin for $C$, by definition it must be a node in $N-F_{C}(N)$, which is a graph that only contains FA nodes of $T$, their support tree descendants in $N$, and possibly transfer nodes that are the parents of FA nodes. Thus either $w$ descends from some $\alpha_{i}\in\alpha_{T}(C)$, or results from a subdivision of $(p_{T}(\alpha_{i}),\alpha_{i})$. In the latter case, $w$ is the parent of $\alpha_{i}$ since an edge can only be subdivided once, as explained above. Therefore, the first condition of our statement holds.
@@
For the second condition, as we now know the origin $w$ descends from some $\alpha_{i}$ or is its parent. Let $\alpha_{j}$ be a FA other than $\alpha_{i}$. Let $l$ be a leaf descending from $\alpha_{j}$ in $\mathcal{T}_{N}$. Since $w$ is an origin, it reaches $l$ in $N-F_{C}(N)$ and thus also in $N$. Since first-appearance nodes are incomparable in $T$ and $\mathcal{T}_{N}$, $w$ is not an ancestor of $l$ in $\mathcal{T}_{N}$. By Lemma 2.2 there is a transfer edge $(u,v)$ in $N$ such that $u\preceq w$ and $l\preceq v$ in $\mathcal{T}_{N}$. The endpoints of this transfer edge must also be in $N-F_{C}(N)$ for $w$ to be able to use it. Next, note that $p_{T}(\alpha_{j})$ has descendants not in $C$. Thus in $N$, $p_{T}(\alpha_{j})$ and its support tree ancestors are in $F_{C}(N)$, and so it follows that $v$ descends from $\alpha_{j}$ or is a transfer node on its parent edge, in which case $v=p_{N}(\alpha_{j})$ (again since edges can only be subdivided once). If $\alpha_{j}$ is a leaf, the latter must hold and we are done. So assume that $\alpha_{j}$ is not a leaf and let $l_{1},l_{2}$ be two leaves descending from $\alpha_{j}$ in $\mathcal{T}_{N}$. Then there are two transfer edges $(u_{1},v_{1})$ and $(u_{2},v_{2})$ that exist in $N-F_{C}(N)$, where $u_{1},u_{2}$ descend from $w$ and $v_{1},v_{2}$ are ancestors of $l_{1}$ and $l_{2}$, respectively. Since these two transfers go out of the $\mathcal{T}_{N}(w)$ subtree, they must in fact be equal by Lemma 2.2. In other words, $w$ reaches every leaf below $\alpha_{j}$ through a single transfer $(u,v)$ such that $v$ is an ancestor of all those leaves. Since $v$ is in $N-F_{C}(N)$, it follows that $v$ must be the parent of $\alpha_{j}$ (and not $\alpha_{j}$ itself because it cannot be a transfer node as its has at least two children, by assumption on $T$). ∎∎
@@
Let $T$ be a galled-completable tree for a character set $\mathcal{C}$. Then for any character $C\in\mathcal{C}$, ${\alpha_{T}(C)\leq 2}$.
@@
Let $N$ be a galled-completion of $T$ that explains $\mathcal{C}$. Suppose for contradiction that $\alpha_{T}(C)\geq 3$ for some $C\in\mathcal{C}$. Let $w$ be an origin for $C$ in $N-F_{C}(N)$. By Lemma 3, $w$ descends from or is the parent of one of the FAs, say $x\in\alpha_{T}(C)$. Let $y,z$ be two other FAs of $C$. Note that, by definition of FAs, $x,y,z$ are pairwise incomparable in $T$. By the second statement of Lemma 3, there are transfer edges $(u_{1},p_{N}(y))$ and $(u_{2},p_{N}(z))$ where $u_{1}$ and $u_{2}$ descend from $w$ in $\mathcal{T}_{N}$. This implies that the subtree $\mathcal{T}_{N}(w)$ contains two outgoing transfers, which contradicts Lemma 2.2. Thus $|\alpha_{T}(C)|\leq 2$ for all $C\in\mathcal{C}$. ∎∎
@@
    If $b_{1}\prec a_{1}$ and $b_{2}$ is not comparable to $a_{1}$, then $a_{2}=b_{2}$.
@@
    If $a_{1}=b_{1}$, then either $b_{2}\prec a_{2}$ or $a_{2}\prec b_{2}$.
@@
For the first statement, if $b_{1}\prec a_{1}$ and $b_{2}$ is outside of the $a_{1}$ subtree, then $b_{1}$ needs to use the same transfer edge as $a_{1}$ to exchange material (because of the galled property). This is only possible if they send the character to the same clade, that is, if $a_{2}=b_{2}$.
For the second statement, if $a_{1}=b_{1}$, then again the two characters must use the same transfer edge to exchange material. In fact, one can argue that $a_{1}=b_{1}$ must be the receiving end of the transfer, and that $a_{2},b_{2}$ must be comparable to be able to use the same node to send the character. ∎
@@
  • Case 1: $a_{2}$ and $b_{2}$ are comparable. Suppose $b_{2}\prec a_{2}$. Let $w$ be an origin for $A$ in $N-F_{A}(N)$. By Lemma 3, $w$ descends from $a_{1}$ or $a_{2}$, or is a transfer node with one of those as a child. Suppose first that $w\preceq a_{1}$ or $w=p_{N}(a_{1})$. We know by Lemma 3 that there exists a transfer edge from some descendant of $w$ that has $p_{N}(a_{2})$ as endpoint. Also by Lemma 3, $p_{N}(b_{2})$ or a descendant of $b_{2}$ is a transfer node whose other endpoint is $p_{N}(b_{1})$ or a descendant of $b_{1}$. This implies that there are two transfer nodes in the subtree $\mathcal{T}_{N}(p_{N}(a_{2}))$ whose other endpoint is outside, which is forbidden by Lemma 2.2. The case where $w\preceq a_{2}$ or is the parent of $a_{2}$ is symmetric.
    Next suppose that $a_{2}\prec b_{2}$. By Lemma 3, one of $p_{N}(a_{1})$ or $p_{N}(a_{2})$ is the receiving end of a transfer, whose sending end is on the other side. If $p_{N}(a_{1})$ is a transfer node, we also know that $p_{N}(b_{1})$ or one of its descendants is a transfer node, with the other endpoint on the $b_{2}$ side. Either way, $\mathcal{T}_{N}(p_{N}(a_{1}))$ has two descending transfer nodes with external endpoints, contradicting Lemma 3. Thus $p_{N}(a_{2})$ is the receiving end of a transfer. By a symmetric argument, $p_{N}(b_{1})$ is the receiving end of a transfer. This implies that there are transfers between $\mathcal{T}_{N}(a_{1})$ and $\mathcal{T}_{N}(b_{2})$ (or their parents) going in opposite directions. Because transfer edges are unidirectional, these transfers are distinct, from which it can easily be seen that there are two underlying cycles intersecting. Therefore, this case is not possible.
  • Case 2: $a_{2}$ and $b_{2}$ are incomparable. By Lemma 3, there is one transfer with one endpoint being $p_{N}(a_{1})$ or a descendant of $a_{1}$, and the other being $p_{N}(a_{2})$ or a descendant of $a_{2}$. Likewise, there is a similar transfer connecting $b_{1}$ and $b_{2}$. Since $a_{2}$ and $b_{2}$ are incomparable, these two transfers must be distinct, they are both in $\mathcal{T}_{N}(a_{1})$ or $\mathcal{T}_{N}(p_{N}(a_{1}))$, and they both have external endpoints since $b_{2}$ is not comparable with $a_{1}$, contradicting Lemma 2.2.
@@
Next, consider the second statement. Let $N$ be a galled-completion of $T$ that explains $\mathcal{C}$. Suppose for a contradiction that $a_{1}=b_{1}$, but neither $a_{2}\prec b_{2}$ nor $b_{2}\prec a_{2}$ holds i.e. $a_{2}$ and $b_{2}$ are incomparable in $\mathcal{T}_{N}$. By Lemma 3, there is a transfer edge between descendants of $a_{1}$ and $a_{2}$ (or their parents), and a transfer edge between descendants of $b_{1}$ and $b_{2}$ (or their parents). Since $a_{2}$ and $b_{2}$ are incomparable, these transfers must be distinct. However, they imply the existence of two transfer nodes descending from $a_{1}=b_{1}$ (or its parent) that have an external endpoint, contradicting Lemma 2.2. ∎∎
Note that in the detailed proof of the first statement of Lemma 3, we use the fact that transfer edges are unidirectional — the statement in the sketch that $b_{1}$ needs to use the “same” transfer as $a_{1}$ remains true, but they do not need to send the character to the same clade in a bidirectional setting, as the sending could go both ways. If we allowed bidirectional transfer edges, we can devise examples in which the lemma does not hold. This shows that even apparently minor changes to the model can lead to more complex structures.
We can begin describing our algorithmic strategy. The first step is to locate and count the FAs for each character to verify the condition established by Lemma 3. A subtree which is rooted at a FA for a specific character $C$ is in fact a maximal clade for $C$. An intuitive way of joining these clades is to add transfer edges between the different subtrees to fulfill the connectivity requirement in Lemma 1. However, this may add superfluous transfer edges, as only the minimal ones are required.
To make this precise, let $v\in V(T)$. We say that another node $w\in V(T)$ is an FA neighbor of $v$ if there exists a character $C\in\mathcal{C}$ such that $\alpha_{T}(C)=\{v,w\}$. We also say that $w$ is a minimal FA neighbor of $v$ if $w\preceq_{T}w^{\prime}$ for every FA neighbor of $v$. A pair of FA nodes $\{v,w\}$ is called simple if $w$ is the unique FA neighbor of $v$ and $v$ is the unique FA neighbor of $w$.
@@
Let $T$ be a tree on leafset $\mathcal{S}$ with character set $\mathcal{C}$. Let $T^{\prime}$ be the tree obtained by subdividing every edge of $T$ once. Then the redundancy-free network for $T$ is the LGT-network $N=(V,E_{S}\cup E_{T})$ with $T^{\prime}$ as support tree obtained as follows:
@@
  • for each $v\in V(T)$ with at least two FA neighbors, and for every minimal FA neighbor $w$ of $v$, add the transfer edge $(p_{T^{\prime}}(w),p_{T^{\prime}}(v))$;
@@
  • for each pair $\{v,w\}$ of simple FA nodes, add one of the transfer edges $(p_{T^{\prime}}(v),p_{T^{\prime}}(w))$ or $(p_{T^{\prime}}(w),p_{T^{\prime}}(v))$ arbitrarily (but not both).
@@
Figure 4: (a) A given tree $T$ on four characters. (b) The FA neighbors between every FAs for each character are indicated by the dashed lines. (c) The redundancy free network of $T$. Note that $C_{3}$ has simple FAs, so the direction for the transfer edge is arbitrary. On the other hand, for $C_{1}$ and $C_{2}$ note that $z$ is the minimal FA neighbor of $y$ so the direction of the transfer edge is not arbitrary.
The idea of this network is that if a node $y$ has multiple FA neighbors, as in Figure 4, then there are many characters with $y$ as a FA. By adding a transfer edge whose sender is the parent of the minimal FA neighbor $z$ of $y$, again as in the figure, all those characters can have their origin on the $z$ side and transfer to $y$. In the figure, we add a transfer from the parent of $z$ to that of $y$ for character $C_{2}$. At the same time, $C_{1}$ can use that same transfer edge because it also wants $y$ and its descendants to inherit the character. For characters whose FAs correspond to a simple pair, no such transfer re-use is needed and the direction is arbitrary.
Before proving our main characterization in terms of redundancy-free networks, we need an intermediary result.

### v3

@@ 2.3 Differences with reconstructions from bipartitions
As we pointed out in the introduction, Warnow et al. [warnow2024statistically] also independently developed galled tree reconstruction algorithms from character data (which are assumed to be SNPs in the paper). With the above definitions in mind, we can now clearly state how PTNs differ from this work. Therein, for each character, the species can be in one of two states, and the states are assumed to correspond to a bipartition of a galled tree $N$. A bipartition is a partition of the leaves that can be obtained by taking a tree displayed by $N$, removing an edge, and taking the two sets of leaves of the resulting connected components (here, “displayed” means a tree that can be obtained by removing one parent edge from each node of in-degree two). They discuss algorithms to reconstruct a galled tree that contains a set of given characters, that is, that contains all the corresponding bipartitions. More specifically, they show that if all the bipartitions of a galled tree $N$ are known (but $N$ is unknown), then $N$ can be reconstructed in polynomial time. They also show that this approach is statistically consistent if characters evolve on a galled tree and may or may not be transmitted on edges labeled as transfer.
@@ 3 The Galled Completion Problem
In other words, $v$ is an FA for $C$ if it roots a maximal subtree of taxa that contain $C$. An example of this definition is shown in Figure 3.
@@
For a character $C\in\mathcal{C}$ and an LGT network $N$, we say that a node $v$ is an origin for $C$ if $v$ reaches every leaf in $C$ in $N-F_{C}(N)$. By Lemma 1, our goal is to add transfer edges to $T$ to ensure that every character has an origin. We first show that in any galled-completion of a tree $T$, an origin of a character $C$ must descend from a FA node (or it could be a transfer node added just above it), and it must “give” its character to all other FAs by transferring just above them.
@@
First note that $V(T)\subseteq V(N)$, by the definition of a completion. For each FA $\alpha\in V(T)$ of character $C$, and each strict ancestor $z$ of $\alpha$ in $T$, i.e., with $\alpha\prec_{T}z$, we have that $L_{T}(z)$ has leaves not containing character $C$. The same is true for $L_{\mathcal{T}_{N}}(z)$, and so $z$ is in $F_{C}(N)$. This means that in $N$, an origin $w$ for $C$ cannot be $p_{T}(v)$ nor an ancestor of $p_{T}(v)$ in $\mathcal{T}_{N}$. So, $w$ must either descend from an FA node $\alpha_{i}$ in $\mathcal{T}_{N}$, or $w$ could be a transfer node inserted above $\alpha_{i}$ while adding transfer edges from $T$ to $N$. For example in Figure 3, in a galled-completion $N$ of $T$, an origin of $C_{2}$ could never be $x$ nor any of its ancestors in $\mathcal{T}_{N}$, because those nodes have $s_{1}$ as a vertical descendant (likewise, $p_{T}(y)$ and its ancestors cannot be origins because of $s_{7}$). This justifies the first part of the statement.
For the second part, if there are multiple FAs, $w$ must reach the ones other than $\alpha_{i}$. Using the definition of FAs, we can deduce that $w$ is incomparable to all nodes of $\alpha_{T}(C)\setminus\{\alpha_{i}\}$ in $\mathcal{T}_{N}$, and Lemma 2 lets us establish that transfer edges as described are needed to achieve this. ∎
@@
We focus on the first condition. Before proceeding, notice that if $N=(V,E_{S}\cup E_{T})$ is a galled-completion of $T$, then $\mathcal{T}_{N}$ is obtained from $T$ by subdividing some edges of $T$ to add transfer nodes, possibly multiple times on the same edge. We claim that $\mathcal{T}_{N}$ can be obtained from $T$ by subdividing each edge at most one time. In other words, no two consecutive transfer nodes can be added along an edge of $T$.
To see this, let $v\in V(T)$ be a non-root node of $T$ and let $p_{v}=p_{T}(v)$ be the parent of $v$ in $T$. Notice that $v,p_{v}\in V(N)$. Suppose for contradiction that from $T$ to $\mathcal{T}_{N}$, the edge $(p_{v},v)$ is subdivided twice or more. That is, suppose that in $N$, there are support edges $(p_{v},x),(x,y)\in E_{S}$, such that $x\neq v$ and $y\neq v$ (and $v\prec_{\mathcal{T}_{N}}y$). Then $x$ and $y$ are both transfer nodes, and are both incident to a distinct transfer edge whose other endpoint is outside of the subtree $\mathcal{T}_{N}(x)$. This contradicts Lemma 2.2 (notice that $x$ and $y$ have a single child in $\mathcal{T}_{N}$ and that the lemma still applies to this case). This lets us establish that some edges of $T$ are subdivided to obtain $\mathcal{T}_{N}$, but each edge can be subdivided at most once.
@@
Let us next observe that if $w$ is an origin for $C$, by definition it must be a node in $N-F_{C}(N)$, which is a graph that only contains FA nodes of $T$, their support tree descendants in $N$, and possibly transfer nodes that are the parents of FA nodes. Thus either $w$ descends from some $\alpha_{i}\in\alpha_{T}(C)$ in $\mathcal{T}_{N}$, or results from a subdivision of $(p_{T}(\alpha_{i}),\alpha_{i})$. In the latter case, $w$ is the parent of $\alpha_{i}$ in $\mathcal{T}_{N}$ since an edge can only be subdivided once, as explained above. Therefore, the first condition of our statement holds.
@@
For the second condition, as we now know the origin $w$ descends from some $\alpha_{i}$ or is its parent in $\mathcal{T}_{N}$. Let $\alpha_{j}$ be an FA other than $\alpha_{i}$. Let $x$ be a leaf descending from $\alpha_{j}$ in $\mathcal{T}_{N}$. Since $w$ is an origin, it reaches $x$ in $N-F_{C}(N)$ and thus also in $N$. Since first-appearance nodes are incomparable in $T$ and $\mathcal{T}_{N}$, $w$ is not an ancestor of $x$ in $\mathcal{T}_{N}$. By Lemma 2 there is a transfer edge $(u,v)$ in $N$ such that $u\preceq w$ and $x\preceq v$ in $\mathcal{T}_{N}$. The endpoints of this transfer edge must also be in $N-F_{C}(N)$ for $w$ to be able to use it. Next, note that $p_{T}(\alpha_{j})$ has descendants not in $C$, both in $T$ and in $\mathcal{T}_{N}$. Thus in $N$, $p_{T}(\alpha_{j})$ and its support tree ancestors are in $F_{C}(N)$, and $v\in V(N-F_{C}(N))$ implies that either $v\prec_{\mathcal{T}_{N}}\alpha_{j}$ or $v$ is a transfer node above $\alpha_{j}$, i.e., $v=p_{N}(\alpha_{j})$ (again since edges can only be subdivided once). If $\alpha_{j}$ is a leaf, the latter must hold and we are done. So assume that $\alpha_{j}$ is not a leaf and let $x_{1},x_{2}$ be two leaves descending from $\alpha_{j}$ in $\mathcal{T}_{N}$. Then there are two transfer edges $(u_{1},v_{1})$ and $(u_{2},v_{2})$ that exist in $N-F_{C}(N)$, where $u_{1},u_{2}$ descend from $w$ in $\mathcal{T}_{N}$ and $v_{1},v_{2}$ are ancestors of $x_{1}$ and $x_{2}$ in $\mathcal{T}_{N}$, respectively. Since these two transfers go out of the $\mathcal{T}_{N}(w)$ subtree, they must in fact be equal by Lemma 2.2. In other words, $w$ reaches every leaf below $\alpha_{j}$ through a single transfer $(u,v)$ such that $v$ is an ancestor of all those leaves. Since $v$ is in $N-F_{C}(N)$, it follows that $v$ must be the parent of $\alpha_{j}$ (and not $\alpha_{j}$ itself because it cannot be a transfer node as its has at least two children, by assumption on $T$). ∎∎
@@
Let $T$ be a galled-completable tree for a character set $\mathcal{C}$. Then for any character $C\in\mathcal{C}$, ${|\alpha_{T}(C)|\leq 2}$.
@@
Let $N$ be a galled-completion of $T$ that explains $\mathcal{C}$. Suppose for contradiction that $|\alpha_{T}(C)|\geq 3$ for some $C\in\mathcal{C}$. Let $w$ be an origin for $C$ in $N-F_{C}(N)$. By Lemma 3, there is $x\in\alpha_{T}(C)$ such that $w\preceq_{\mathcal{T}_{N}}x$ or $w=p_{N}(x)$. Let $y,z$ be two other FAs of $C$. Note that, by definition of FAs, $x,y,z$ are pairwise incomparable in $T$, and thus in $\mathcal{T}_{N}$. By the second statement of Lemma 3, there are transfer edges $(u_{1},p_{N}(y))$ and $(u_{2},p_{N}(z))$ in $N$ where $u_{1}$ and $u_{2}$ descend from $w$ in $\mathcal{T}_{N}$. This implies that the subtree $\mathcal{T}_{N}(w)$ contains two outgoing transfers, which contradicts Lemma 2.2. Thus $|\alpha_{T}(C)|\leq 2$ for all $C\in\mathcal{C}$. ∎∎
@@
    If $b_{1}\prec_{T}a_{1}$ and $b_{2}$ is not comparable to $a_{1}$ in $T$, then $a_{2}=b_{2}$.
@@
    If $a_{1}=b_{1}$, then either $b_{2}\prec_{T}a_{2}$ or $a_{2}\prec_{T}b_{2}$.
@@
For the first statement, by Lemma 3 in any galled-completion $N$ of $T$ there is a transfer edge with one end that is either $p_{N}(a_{1})$ or descends from $a_{1}$ in $\mathcal{T}_{N}$, and the other end is outside of $\mathcal{T}_{N}(a_{1})$. Likewise, some transfer edge has one end that is $p_{N}(b_{1})$ or descends from $b_{1}$, and the other end is $p_{N}(b_{2})$ or it descends from $b_{2}$. That other end is also outside of $\mathcal{T}_{N}(a_{1})$, since $a_{1}$ and $b_{2}$ are incomparable (in both $T$ and $\mathcal{T}_{N}$) by assumption. Hence by Lemma 2.2 the two transfer nodes must be equal, i.e., $A$ and $B$ use the same transfer edge in $N$ to transfer the character. With some case analysis, we can show that this is only possible if the origins of both $A$ and $B$ are on the $a_{1}$ and $b_{1}$ side, and they send their character to $a_{2}$ and $b_{2}$ using that transfer edge. The second part of Lemma 3 lets us deduce that the receiving end must be $p_{N}(a_{2})=p_{N}(b_{2})$ and that $a_{2}=b_{2}$.
For the second statement, if $a_{1}=b_{1}$, then again the two characters must use the same transfer edge to exchange material. In fact, one can argue that $p_{N}(a_{1})=p_{N}(b_{1})$ must be the receiving end of the transfer, and that $a_{2},b_{2}$ must be comparable to be able to use the same transfer edge to send the character. ∎
@@
  • Case 1: $a_{2}$ and $b_{2}$ are comparable. Suppose $b_{2}\prec_{T}a_{2}$. Let $w$ be an origin for $A$ in $N-F_{A}(N)$. By Lemma 3, $w$ descends from $a_{1}$ or $a_{2}$ in $\mathcal{T}_{N}$, or is a transfer node with one of those as a child. Suppose first that $w\preceq_{\mathcal{T}_{N}}a_{1}$ or $w=p_{N}(a_{1})$. We know by Lemma 3 that there exists a transfer edge from some descendant of $w$ in $\mathcal{T}_{N}$ that has $p_{N}(a_{2})$ as endpoint. Also by Lemma 3, $p_{N}(b_{2})$ or a descendant of $b_{2}$ in $\mathcal{T}_{N}$ is a transfer node whose other endpoint is $p_{N}(b_{1})$ or a descendant of $b_{1}$. But $b_{2}\prec_{\mathcal{T}_{N}}a_{2}$ implies that these two transfer edges are distinct, and so there are two transfer nodes in the subtree $\mathcal{T}_{N}(p_{N}(a_{2}))$ whose other endpoint is outside, which is forbidden by Lemma 2.2. The case where $w\preceq_{\mathcal{T}_{N}}a_{2}$ or is the parent of $a_{2}$ is symmetric.
    Next suppose that $a_{2}\prec_{\mathcal{T}_{N}}b_{2}$. By Lemma 3, one of $p_{N}(a_{1})$ or $p_{N}(a_{2})$ is the receiving end of a transfer, whose sending end is on the other side. If $p_{N}(a_{1})$ is a transfer node, we also know that $p_{N}(b_{1})$ or one of its descendants in $\mathcal{T}_{N}$ is a transfer node, with the other endpoint on the $b_{2}$ side. Therefore, $\mathcal{T}_{N}(p_{N}(a_{1}))$ has two descending transfer nodes with external endpoints, contradicting Lemma 3. Thus $p_{N}(a_{2})$ is the receiving end of a transfer. By a symmetric argument, $p_{N}(b_{1})$ is the receiving end of a transfer. This implies that there are transfers edges between nodes of $\mathcal{T}_{N}(a_{1})$ and nodes of $\mathcal{T}_{N}(b_{2})$ (or their parents) going in opposite directions. Because transfer edges are unidirectional, these transfers are distinct, from which it can easily be seen that there are two underlying cycles intersecting. Therefore, this case is not possible.
  • Case 2: $a_{2}$ and $b_{2}$ are incomparable. By Lemma 3, there is one transfer edge with one endpoint being $p_{N}(a_{1})$ or a descendant of $a_{1}$ in $\mathcal{T}_{N}$, and the other being $p_{N}(a_{2})$ or a descendant of $a_{2}$ in $\mathcal{T}_{N}$. Likewise, there is a similar transfer edge connecting $b_{1}$ and $b_{2}$. Since $a_{2}$ and $b_{2}$ are incomparable, these two transfers edges must be distinct. Because $b_{1}\prec_{\mathcal{T}_{N}}a_{1}$, they are both in $\mathcal{T}_{N}(a_{1})$ or $\mathcal{T}_{N}(p_{N}(a_{1}))$, and they both have an outside endpoint since $b_{2}$ is not comparable with $a_{1}$, contradicting Lemma 2.2.
@@
Next, consider the second statement. Let $N$ be a galled-completion of $T$ that explains $\mathcal{C}$. Suppose for a contradiction that $a_{1}=b_{1}$, but neither $a_{2}\prec_{T}b_{2}$ nor $b_{2}\prec_{T}a_{2}$ holds, i.e., $a_{2}$ and $b_{2}$ are incomparable in $\mathcal{T}_{N}$. By Lemma 3, there is a transfer edge between descendants of $a_{1}$ and $a_{2}$ (or their parents) in $\mathcal{T}_{N}$, and a transfer edge between descendants of $b_{1}$ and $b_{2}$ (or their parents) in $\mathcal{T}_{N}$. Since $a_{2}$ and $b_{2}$ are incomparable, these transfers must be distinct. However, they imply the existence of two transfer nodes descending from $a_{1}=b_{1}$ (or its parent) in $\mathcal{T}_{N}$ that have an external endpoint, contradicting Lemma 2.2. ∎∎
Note that in the detailed proof of the first statement of Lemma 3, we use the fact that transfer edges are unidirectional — the statement in the sketch that $A$ and $B$ need to use the “same” transfer edge remains true, but they do not need to send the character in the same direction in a bidirectional setting, as the sending could go both ways. If we allowed bidirectional transfer edges, we can devise examples in which the lemma does not hold. This shows that even apparently minor changes to the model can lead to more complex structures.
An algorithm using redundancy-free networks
We can begin describing our algorithmic strategy. The first step is to locate and count the FAs for each character to verify the condition established by Lemma 3. A subtree which is rooted at an FA for a specific character $C$ is in fact a maximal clade for $C$. An intuitive way of joining these clades is to add transfer edges between the different subtrees to fulfill the connectivity requirement in Lemma 1. However, this may add superfluous transfer edges, as only the minimal ones are required.
To make this precise, let $v\in V(T)$. We say that another node $w\in V(T)$ is an FA neighbor of $v$ if there exists a character $C\in\mathcal{C}$ such that $\alpha_{T}(C)=\{v,w\}$. We also say that $w$ is a minimal FA neighbor of $v$ if $w\preceq_{T}w^{\prime}$ for every FA neighbor of $v$. A pair of FA nodes $\{v,w\}$ is called simple if $w$ is the unique FA neighbor of $v$ and $v$ is the unique FA neighbor of $w$. See Figure 4.b for an example, where FA neighborhood relationships are shown in dotted lines.
It turns out that these relationships tell us where transfer edges should be added.
@@
Let $T$ be a tree on leafset $\mathcal{S}$ with character set $\mathcal{C}$. Let $T^{\prime}$ be the tree obtained by subdividing every edge of $T$ once. Then a redundancy-free network for $T$ is an LGT-network $N=(V,E_{S}\cup E_{T})$ with $T^{\prime}$ as support tree obtained as follows:
@@
  • for each $v\in V(T)$ with at least two FA neighbors, and for every minimal FA neighbor $w$ of $v$, add the transfer edge $(p_{T^{\prime}}(w),p_{T^{\prime}}(v))$ to $E_{T}$;
@@
  • for each pair $\{v,w\}$ of simple FA nodes, add to $E_{T}$ one of the transfer edges $(p_{T^{\prime}}(v),p_{T^{\prime}}(w))$ or $(p_{T^{\prime}}(w),p_{T^{\prime}}(v))$ arbitrarily (but not both).
Note that there may be multiple redundancy-free networks $N$ for a tree $T$, since the direction of transfer edges added in the second step is arbitrary. Let us also point out that if $w$ is an FA neighbor of $v$, then $v,w$ are FAs of some character $C$ and are therefore incomparable. Hence, edges of $N$ are only between incomparable nodes.
@@
Figure 4: (a) A given tree $T$ on four characters. (b) The FA neighbors between every FAs for each character are indicated by the dashed lines. In this example, $y$ is an FA for $C_{1}$ and $C_{2}$ and has two FA neighbors $x$ and $z$ ($x$ is an FA for $C_{1}$ and $z$ an FA for $C_{2}$). Here $z$ is a minimal FA neighbor of $y$ because no strict descendant of $z$ is an FA neighbor of $y$, but $x$ is not a minimal FA neighbor of $y$ (because $z\prec_{T}x$). (c) A redundancy free network of $T$. Note that $C_{3}$ has simple FAs, so the direction for the transfer edge is arbitrary. On the other hand, for $C_{1}$ and $C_{2}$ note that $z$ is the minimal FA neighbor of $y$ so the direction of the transfer edge is not arbitrary.
To illustrate the idea of a redundancy-free network and the importance of minimal FA neighbors, consider the example provided in Figure 4. Therein, node $y$ is an FA for two characters $C_{1},C_{2}$ and has two FA neighbors $x$ and $z$, which are respectively FAs of $C_{1}$ and $C_{2}$. Here, $z$ is a minimal FA neighbor of $y$ but not $x$, as seen on subfigure (b). By Lemma 3, in any galled-completion $N$ of $T$, because of $C_{1}$ either $p_{N}(x)$ or a descendant of $x$ in $\mathcal{T}_{N}$ is a transfer node, and because of $C_{2}$ we must have that $p_{N}(z)$ is a transfer node (because in the example $z$ is a leaf and cannot have descending transfer nodes). By the same lemma, both these transfer nodes are incident to an edge whose other end is $p_{N}(y)$ or a descendant of $y$, and to satisfy Lemma 2.2 these two transfer edges must be equal (otherwise we create intersecting cycles). In other words, $C_{1}$ and $C_{2}$ must use the same transfer edge to connect their FAs, and one end of that transfer edge must be the parent of $z$, the minimal FA neighbor of $y$. One can then work out that the edge must be from $p_{N}(z)$ to $p_{N}(y)$, since otherwise $x$ has no way to receive the character $C_{1}$ from a transfer to transmit it vertically. This is precisely the edge that we add to the redundancy-free network.
This idea generalizes as follows. Suppose that some node $y$ of $T$ has multiple FA neighbors $x_{1}\prec_{T}x_{2}\prec_{T}\ldots\prec_{T}x_{k}$, where $x_{1}$ is a minimal FA neighbor of $y$, because of some character $C$. As argued above, for each $i\in[k]$ some transfer edge is needed between $p_{N}(x_{i})$ or a descendant of $x_{i}$, and $p_{N}(y)$ or a descendant of $y$. Still by Lemma 2.2, all these transfer edges must be equal, and one end of that transfer must be $p_{N}(x_{1})$ or a descendant of $x_{1}$. Hence, the minimal FA neighbor $x_{1}$ of $y$ gives the “highest” possible location at which a transfer node can be added. Then, the direction from the $x_{1}$ side to the $y$ side is forced since all the $x_{i}$’s must use that transfer edge to send their character to all descendants of $y$. We add such an edge in the construction of a redundancy-free network.
Before proving our main characterization in terms of redundancy-free networks, we need an intermediary result, which essentially states that if $T$ satisfies all the established properties so far, then each node of $N$ needs to send its characters to at most one other node. In particular, we never add bidirectional edges, that is, for two nodes $u,v$ at most one of $(u,v)$ or $(v,u)$ is added.